Ask
your own question, for FREE!
Mathematics
9 Online
A bowl contains candies of the same size in three flavors: orange, strawberry, and pineapple. If the probability of randomly pulling out an orange candy is fraction 2 over 8, and the probability of randomly pulling out a strawberry candy is fraction 3 over 8, what is the probability of randomly pulling out a pineapple candy?
Still Need Help?
Join the QuestionCove community and study together with friends!
2/8+3/8 =5/8?
2/8-3/8 = 1/8?
64% is what i got and ill show you how...
ok
We need to end up with \(\Large\rm \frac{8}{8}\) with our three individual odds.
Still Need Help?
Join the QuestionCove community and study together with friends!
So far we have \(\Large\rm \frac{2}{8}\) and \(\Large\rm \frac{3}{8}\)
o then the pineapple would be 3/8
since they should be equal?
so im sorry i meant 3/8 not 64%-i dont know where that came from :b
yes! :)
Still Need Help?
Join the QuestionCove community and study together with friends!
thanks @Kim21 got it
welcome :)
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
8 hours ago
1 Reply
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal